NaOH H ? 1 eg 2 eq Ph Ph. Ph Ph Ph Ph Ph O A B С. D

Is this right?
Answer #1

CH3-E - CHz on CH3 C-CHZ 25 ul CH3-C - CH2 + H G ph H20 [ proton transfery OH u CH3-C-CH2-6-ph H A t-H2O O CH3 -C- CH = CH -

AL-H₂O C) =cH- Ph CH3 -C-CH C–CH = CHỮ оде CH2 -CH=CH-ph с – сHз Ph-CH=CH-C-CH₂ ДНЗ » 0 ~Нt++ T на нема | | -CH3 Нас | Ph be

The reaction is an example of crossed aldol reaction.

so,  correct option is B.

Your answer is right.